OpenStudy (anonymous):

Solve this system of equations. Please write your answer as an ordered pair. x+3y=1 -3x-3y=-15

OpenStudy (anonymous):

this time lets set it equal to x x+3y=1 in slope intercept form it is x=-3y+1 plug in x=-3y+1 to -3x-3y=-15 -3(-3+1)-3y=-15 9-3=6-3y=-15=y=7

OpenStudy (anonymous):

plug in y to any of the equations to get x

OpenStudy (anonymous):

first equation is easier so x+3(7)=1=x+21=1 x=-20 answer is (-20,7)

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

more...